OpenStudy (anonymous):

Given the function f(x) = 2^x, find the value of f^−1(16).

ganeshie8 (ganeshie8):

set f(x) = 2^x = 16, and solve x

OpenStudy (anonymous):

what will i do with the f^--1(16)

ganeshie8 (ganeshie8):

whatever value u get for x, when u solve below :- 2^x = 16 is the value of f^-1(16)